problem: solve for w: 7 (is less then /equal to 5 minus 2w is less then/ equal to 14
Answer by tutorcecilia(2152)

7<=5-2w<=14 [Subtract -5 from all three sides of the inequality]
7-5<=5-5-2w<=14-5 [Simplify]
2<=-2w<=9
2/2<=-2w/2<=9/2[Divide by 2 on all three sides of the inequality]
1<=(-1)w<=9/2 [Simplify]
1/-1>=(-1)w/-1>=(9/2)\(-1)[Divide all three sides by -1 and switch the signs]
-1>=w>=-9/2 [Final answer][When dividing or multiplying by a negative number, switch the inequality signs]
.
.
Check the final answer to see if it makes sense.
Plug-in any number for "w" as long as -1 is greater than or equal to that number "w", yet the number "w" is greater than or equal to -4.5.
Use the number line as your guide.
We can pick numbers such as in -2, -3 or -4.
Let's try -2. Go back to the original equation and plug it in.
7<=5-2(-2)<=14
7<=9<=14 [Checks out. This is a true statement]
Please let me know if this helps. Thanks.